Shiftsum for DAY shift 10/Jan/2000


Subject: Shiftsum for DAY shift 10/Jan/2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:15:58 CST


Shift summary for day shift 01/10/2000

On Shift: AT, RFB

Hours of beam available at all intensities: 5.4
Hours of beam available at nominal intensity: 5.4
Hours of E799 data (special runs not included): 2.8
Hours of special runs (laser scans, muon runs...): 0.75

Accounting of lost time:

A TeV magnet quench. Starting and Stopping runs, KFREND's. Changed power
supply at crate 5.
========================================================================

Tapes brought to FCC this shift:

None
========================================================================

Accesses during shift: Two

Bob, Jason LaDue and Hogan swapped a power supply at crate 5.
Jason Hamm and Arun to reset Et power suplly.

========================================================================

Runs ended this shift:

15456 - 0.2M - KAMI_KE3
15457 - - Ktev Ped
15458 - 0.5M - Muon Run
15459 - 0.6M - BEAM_799
15460 - 2.7M - BEAM_799

=======================================================================

Overview of shift:

KAMI_KE3 data-taking, E799 data-taking, and some problems.

========================================================================

Description of Shift:

run spill time comment
------------------------------------------------------------------------

15445 08:20 Rebooted ATEV3. Jason Hamm working with Etotal
                       power state.

                08:50 Restored BA HV for E799.

                08:37 Etotal error, hcc error and trigger
                       hangs. Reload ETOT maps + a number of KFREND's.
                       ET had tripped. Jason came in, and reset the
                       power supply.

                09:00 Problems with LASER ILLUMINATION: Chn 496.

15457 10:02 Start a Ped run.

                10:08 End of Ped run.

                10:10 CsI Channels 490-520 had a HV glitch. Update
                       ADGN constants for run 15457.

15458 10:15 Start a Muon Run.

                10:30 End of Muon Run.

15458 10:40 Start new BEAM_799 run 15459.

                11:00 Pipeline errors. Trying to rebooted pipeline
without
                       success. Paging BobT. There is a blown pipeline
power supply.
                       BobT and jason Ladue swapped it.

                11:40 Made access to check voltages or RR51 - crate,
which
                       turned out to be OK.

                12:10 End of run 15459.

15460 12:17 Started new BEAM_799 run 15460.

                13:42 TeV Ramp problem.

                14:27 Power Supply problem.

                14:33 Beam back!!!

                14:40 Tried to put crate 16 DYC 16 into readout. Got
full
                       at high rate. Take it out. Update ADGN constants
from run 15460. Stop run 15460.

                14:56 Damp didn't start; logout and drt_fresh_start

                15:25 Failed again to start a new run. Probably drive 4
at
                       ktev2 is dead. Finnaly it was a bad tape. Changed
the tape and started the run 15466.

========================================================================
Items to Follow up from previous Shift

  Next run do DRT_FRESH _START
  
  Label 30 QKE tapes for short reg run and other special runs
 
  Don't forget to hit "reset quartz" button after each KFREND.

========================================================================
Items to Follow up on Next Shift

The next shift is alread here!!! Just pray for a good shift!!!!

-- 
+---------------------------------------+
|					|
|	Ricardo Franco Barbosa		|
|		  diz			|
|		Ate' Logo!		|
|					|
+---------------------------------------+



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:15:59 CST